95c:20072 20J05 18G60 55U99

Mislin, Guido ^{(}CH-ETHZ^{)}

Tate cohomology for arbitrary groups via satellites. (English. English summary) Topology Appl. 56 (1994), no. 3, 293–300. Summary: “We define cohomology groups Ĥ^{n}(G; M), n ∈ Z, for an arbitrary group G and a G-module M, using the concept of satellites. These cohomology groups generalize the Farrell-Tate groups for groups of finite virtual cohomological dimension and form a connected sequence of functors, characterized by a natural universal property.”

Yining Xia (1-NIL)

\begin{matrix}c \\ ◯ \end{matrix} Copyright American Mathematical Society 1995, 1996